What are magnitude and direction [1 pt.] of the m
Answer #1

Given

Current I = 1.50 A

Inner radius of quadrant toroid a = 5.0 cm = 0.05 m

Outer radius of quadrant toroid a = 8.0 cm = 0.08 m

length of the quadrant toroid L = 2\pi(a + b)/4 + 2(b - a) = 2\pi(0.05 + 0.08)/4 + 2(0.08 - 0.05) = 0.2642 m

Now according to toroid, the magnetic field is given by

B = \muoI/L = (4\pix10-7x1.5x8/0.2642) = 7.1x10-6 Tesla